OpenStudy (anonymous):

What is the standard form equation of the line passing through (5, –3) with a slope of –3?

OpenStudy (anonymous):

appy y-y1=m(x-x1)

OpenStudy (hoblos):

y+3= -3(x-5) y=-3x +12 3x+y -12=0

OpenStudy (anonymous):

3x – y = 12 3x – y = –12 3x + y = 12 3x + y = –12